write the equation of the line that contains the
indicated point(s), and/or has the given slope or intercepts; use
either the slope–intercept form y = mx + b, or the form x = c.
(4, 6); (4, -3)

Answer by TimothyLamb(4379)

slope = m = dy/dx
m = (6+3)/(4-4)
m = (9)/(0)
division by zero is undefined:
slope is undefined:
vertical line:
x = 4
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---